# Quantum Machine Learning: from NISQ to Fault Tolerance

### v1

@@ III.3 Harrow-Hassidim-Lloyd algorithm
In this section, we will delve into the intricacies of the HHL algorithm, as detailed in Harrow et al. (2009). The fundamental setup remains consistent:
$A\ket{x}=\ket{b}~{},$ (12)
@@
|$\ket{x}=A^{-1}\ket{b}=\sum_{i,j=0}^{N-1}\lambda_{i}^{-1}|u_{i}\rangle\langle u_{i}|b_{j}\ket{u_{j}}=\sum_{i,j=0}^{N-1}\lambda^{-1}_{i}b_{j}\delta_{ij}\ket{u_{i}}=\sum_{i=0}^{N-1}\lambda^{-1}_{i}b_{i}\ket{u_{i}}~{}.$ (13)
@@
Figure 6: Circuits for HHL algorithm au2 et al. (2023).
@@
where constant $C$ is the normalization factor. If we obtain |$\ket{0}~{},$ hen we will have to repeat the above procedure until we obtain the desired result. Now we have our desired $\lambda_{j}^{-1}$, so we can apply the inverse of the quantum phase estimation procedure and return to our initialized second register and we have the following identification:
@@
The state output of the first register is our desired solution state or vector. The whole post-selection process is for us to be sure that the state in the first register is now proportional to the solution vector, and from there, we can complete our algorithm. For a more detailed discussion of the HHL algorithm, please refer to the original paper Harrow et al. (2009). The HHL algorithm takes $\mathcal{O}((\log{N})^{2})$ quantum steps to output |$\ket{x}$, compared with $\mathcal{O}(N\log{N})$ steps required to find the vector $\bm{x}$ using the best-known method on a classical computer. HHL does achieve a significant speedup, and so does the quantum SVM compared to its classical counterpart.
Before we become overly enthusiastic about this algorithm and its broad applications, it’s important to note the following caveats, which can be crucial in practice Aaronson (2015).
@@
    $\bm{b}\rightarrow\ket{b}=\sum_{i=0}^{N-1}b_{i}\ket{u_{i}}~{},$ (16)
    must occur rapidly. If the preparation of the state |$\ket{b}$ requires exponential steps, the significant speedup gained by HHL in subsequent steps might be negated. Theoretically, this might be achieved through quantum random access memory (QRAM), which will be discussed in section III.4. Nevertheless, it is crucial that $\bm{b}$ is reasonably uniform. For evident reasons, this challenge is referred to as the input problem Aaronson (2015); Biamonte et al. (2017).
@@
    Finally, there’s the output problem, which is the converse of the input problem Biamonte et al. (2017). The challenge of encoding classical information into quantum states also extends to reading the output quantum state |$\ket{x}$ into classical information $\bm{x}$ when HHL is finished. Learning the value of any specific entry $x_{i}$ by measurement will, in general, require repeating the algorithm roughly $N$ times, negating the significant speedup.
To summarize, HHL is an algorithm designed to solve a system of linear equations approximately in logarithmic time. It serves as a template for other quantum algorithms, as long as one can address and mitigate the various caveats associated with its practical implementation Aaronson (2015). If these challenges are properly handled, HHL can find applications in various real-world scenarios.
III.3.1 Large scale machine learning models in fault tolerant era
@@
The paper Liu et al. (2021) introduces a technique known as Carleman linearization, demonstrating that non-linear ODEs can be linearized into an infinite set of linear ODEs that can be truncated to a chosen level of approximation. This allows the approximate solution of challenging non-linear ODEs by solving the easier linearized versions. The authors establish that if the original ODEs are non-linear and dissipative, then the error resulting from the truncation in Carleman linearization can be controlled. This implies that a non-linear, dissipative system can be efficiently linearized.
@@
Dissipativeness plays a crucial role in various machine learning tasks. Many machine learning processes, including gradient descent, exhibit dissipative behavior. Treating gradient descent as dynamics reveals that it operates as an open first-order system outside the scope of Lagrangian mechanics Liu et al. (2023). Biological processes in the brain also exemplify the significance of dissipation, where forgetting becomes essential for making room to remember new information Liu et al. (2023).
The findings of Liu et al. (2023) suggest the potential for an efficient algorithm to solve stochastic gradient descent. Generalizing the techniques for differential equations to ordinary difference equations, it is observed that in the learning process, the Hessian eigenvalues characterize the degree of dissipation. Towards the end of training, there are roughly equal numbers of positive and negative eigenvalues, signifying comparable dissipative and non-dissipative modes. However, at the beginning of training, the number of dissipative modes significantly exceeds the number of non-dissipative modes, creating highly asymmetric Hessian eigenvalues. This phase is recognized as highly dissipative, consistent with the intuition that the system learns rapidly in the initial stages. The ODE solver described earlier is applicable during this phase, enabling the solution of machine learning tasks. Towards the end of training, when positive and negative eigenvalues become comparable, the algorithm may not be suitable for solving ODEs in general. An exception is when approaching a local minimum, where the Hessian matrix is positive definite, indicating a dissipative system suitable for the ODE solver. In summary, this quantum algorithm might help machine learning processes significantly, and if the Hessian becomes symmetric towards the end, classical algorithms can be employed for further computation.
@@
Indeed, the vision of leveraging quantum algorithms, particularly with large-scale implementations of algorithms like the HHL, holds the promise of addressing the increasing costs associated with classical algorithms. The potential gains in efficiency and computational speed offered by quantum computers could revolutionize machine learning and other computational tasks. However, it’s important to emphasize that realizing these advancements requires the implementation of full quantum error correction to ensure fault tolerance. Overcoming the challenges associated with error correction remains a critical aspect of bringing quantum computing applications to fruition. Moreover, some further explorations about machine learning foundations should be given along this line, like improving the algorithms from solving gradient descent to solving the matrix multiplication problems in large-scale classical neural networks for back propagation, where quantum linear algebra techniques or parallel computing from analog photonic devices might be helpful for those tasks Becker et al. (2023).
III.4 Quantum random access memory
@@
Now, let’s explore the challenges associated with the “input problem” and the “output problem”. Classical data needs to be inputted into a quantum computer before undergoing processing Biamonte et al. (2017). Referred to as the ‘input problem’, this step is generally performed with minimal overhead but can pose a significant bottleneck for certain algorithms. Similarly, the ‘output problem’ arises when retrieving data after it has been processed on a quantum device. Much like the input problem, the output problem can lead to a substantial operational slowdown.
@@
In particular, when considering the application of algorithms like HHL, least squares fitting, quantum support vector machines, and related approaches, substantial amounts of classical data might be loaded into a quantum system. This loading process can demand significant time, as acknowledged in the literature Aaronson (2015); Biamonte et al. (2017). While the use of a quantum random access memory (QRAM) Hann (2021); Hann et al. (2019a) might, in principle, address this issue, the associated costs might be prohibitive for large-scale data problems Arunachalam et al. (2015).
@@
QRAM can be conceptualized as an architecture designed for the implementation of quantum oracles Hann (2021). Consider a computational problem where the input is represented by a classical data vector $\bm{x}~{}.$ An oracle, often described as a black box, is a tool that can be interrogated to disclose information about $\bm{x}~{}.$ Although the oracle reveals details about $\bm{x}$ upon request, the specific method it employs to obtain this information is not explicitly specified.
Figure 7: Classical and quantum data-lookup oracles Hann (2021).
@@
Oracles can exist in either classical or quantum forms, as illustrated in FIG. 7. A basic example of a classical oracle is the data-lookup oracle. This type of oracle is activated by providing it with an index $i$ as input, and it subsequently outputs the corresponding vector element $x_{i}~{}.$ A natural extension of this classical data-lookup oracle is the quantum data-lookup oracle. In the quantum scenario, both the inputs and outputs of the oracle query are represented as quantum states. The query itself is implemented by a unitary operation, denoted as $\mathcal{O}^{(\text{DL})}_{x}$, which performs the mapping Hann (2021):
@@
where the notation $b$ represents any computational basis state and $\oplus$ denotes addition modulo 2. The superscripts $A$ and $B$ refer to two quantum registers, the state of register $A$ indicates which element to look up, and the query encodes this element into the state of register $B~{}.$ It’s important to note that, while the inputs and outputs of the query are quantum, the data being queried is classical. This characteristic makes quantum oracles serve as a bridge between classical data and quantum algorithms Hann (2021).
However, merely loading classical data into quantum memory is not sufficient; the process needs to be fast Aaronson (2015). QRAM is designed to address this requirement. In QRAM, the quantum state summarizing the vector uses $\log{d}$ qubits, and the QRAM operation involves $\mathcal{O}(d)$ operations distributed over $\mathcal{O}(\log{d})$ steps, which can be performed in parallel Biamonte et al. (2017).
For QRAM, a quantum superposition of different addresses |$\ket{\psi_{\text{in}}}$ serves as input, and the QRAM produces an entangled state |$\ket{\psi_{\text{out}}}$ where each address is correlated with the corresponding memory element Hann (2021):
@@
Figure 8: The left figure is a picturesque illustration of fanout QRAM Hann (2021) to retrieve information stored at $i=5~{}.$ On the other hand, the right figure is a picturesque illustration of bucket-brigade QRAM Hann (2021) to retrieve information stored at $i=5~{}.$
@@
The bucket-brigade QRAM is a bit different in the first stage. We don’t flip all the router qubits in the layer corresponding to the address qubits. Instead, we route the address qubits into the QRAM one by one, with the router qubits initialized at some state |$\ket{w}~{}.$ Since we route in the address qubits in order, we naturally carve out the path that will lead to the information stored in $x_{5}~{}.$
@@
There are debates about whether QRAM is actually practical since it is time-saving but also hardware-consuming. Recently, there are several different QRAM designs that are hardware-efficient. One example is developed by the authors of Hann et al. (2019b). They made use of circuit quantum acoustodynamic (cQAD) systems O’Connell et al. (2010); Gustafsson et al. (2014) and constructed quantum gates by applying external drives with different frequencies.
@@
Furthermore, due to its logarithmic depth nature, when QRAM is stored with a large number of data, the speed for querying it will ultimately be bounded by the speed of light. The problem is how many qubits QRAM can store before we hit the boundary set by locality. This problem is addressed in Wang et al. (2023), where the authors considered locality and found that some 2D or 3D QRAM designs can handle up to $10^{20}$ qubits before hitting the constraint line. This might be already a sufficient number of qubits to perform machine learning training programs. As an result, QRAM might be still reasonable to build in the future.
@@
Although it looks like that QRAM might be implementable, it is also important that there is no definitive evidence or proof in the hardware especially in the large scale up to date. So whether we can achieve exponential speed ups that some quantum machine learning programs promises might be still unknown and controversial. On the other hand, for some of the QML programs, one can go around QRAM by implementing a quantum subroutine which prepares the quantum states that encodes classical information Biamonte et al. (2017).
IV Statistical learning theory
@@
In this section, our aim is to approach a quantum counterpart to statistical learning theory, initiating a systematic framework for delving into the fundamentals of quantum machine learning. A crucial aspect intertwined with statistical learning theory is shadow tomography, addressing a challenge rooted in the distinctive nature of quantum measurements—destructiveness. When concluding a Quantum Machine Learning (QML) program and obtaining a quantum state, extracting classical information becomes a non-trivial endeavor. The challenge arises from the inevitable collapse of the state during measurement, limiting access to the original state. Working with a single copy of an unknown quantum state, denoted as $\rho$, poses an inherent difficulty, as clever strategies cannot approximate a classical description of $\rho$ through direct measurement. The overarching task of reconstructing a description of a $D$-dimensional quantum mixed state $\rho$,” given multiple copies of $\rho$, falls under the domain of quantum state tomography Aaronson (2018).
@@
One significant development inspired by shadow tomography is known as classical shadow. As discussed earlier in section II.1.3, this approach holds promise for addressing the output or downloading problem. Classical shadow infers the output state $\rho$ by performing random measurements, averaging the obtained classical data, and then inverting the results to retrieve the quantum state. This method extensively leverages basics of probability theory to control errors, making it an efficient quantum-to-classical data converter Elben et al. (2022). Utilizing classical shadow, one can devise classical machine learning algorithms that exhibit only polynomial differences with full fault-tolerant QML programs, particularly when focusing on average error Huang et al. (2021a). Moreover, this approach holds significant applications in QML for quantum data or quantum simulators, especially when a readout is required at the end.
IV.1 Shadow tomography

### v2

@@ 3.3 Harrow-Hassidim-Lloyd algorithm
In this section, we will delve into the intricacies of the HHL algorithm, as detailed in [58]. The fundamental setup remains consistent:
$A\ket{x}=\ket{b}\leavevmode\nobreak\ ,$ (12)
@@
|$\ket{x}=A^{-1}\ket{b}=\sum_{i,j=0}^{N-1}\lambda_{i}^{-1}|u_{i}\rangle\langle u_{i}|b_{j}\ket{u_{j}}=\sum_{i,j=0}^{N-1}\lambda^{-1}_{i}b_{j}\delta_{ij}\ket{u_{i}}=\sum_{i=0}^{N-1}\lambda^{-1}_{i}b_{i}\ket{u_{i}}\leavevmode\nobreak\ .$ (13)
@@
Figure 6: Circuits for HHL algorithm [11].
@@
where constant $C$ is the normalization factor. If we obtain |$\ket{0}\leavevmode\nobreak\ ,$ hen we will have to repeat the above procedure until we obtain the desired result. Now we have our desired $\lambda_{j}^{-1}$, so we can apply the inverse of the quantum phase estimation procedure and return to our initialized second register and we have the following identification:
@@
The state output of the first register is our desired solution state or vector. The whole post-selection process is for us to be sure that the state in the first register is now proportional to the solution vector, and from there, we can complete our algorithm. For a more detailed discussion of the HHL algorithm, please refer to the original paper [58]. The HHL algorithm takes $\mathcal{O}((\log{N})^{2})$ quantum steps to output |$\ket{x}$, compared with $\mathcal{O}(N\log{N})$ steps required to find the vector $\bm{x}$ using the best-known method on a classical computer. HHL does achieve a significant speedup, and so does the quantum SVM compared to its classical counterpart.
Before we become overly enthusiastic about this algorithm and its broad applications, it’s important to note the following caveats, which can be crucial in practice [1].
@@
    $\bm{b}\rightarrow\ket{b}=\sum_{i=0}^{N-1}b_{i}\ket{u_{i}}\leavevmode\nobreak\ ,$ (16)
    must occur rapidly. If the preparation of the state |$\ket{b}$ requires exponential steps, the significant speedup gained by HHL in subsequent steps might be negated. Theoretically, this might be achieved through quantum random access memory (QRAM), which will be discussed in section 3.4. Nevertheless, it is crucial that $\bm{b}$ is reasonably uniform. For evident reasons, this challenge is referred to as the input problem [1, 18].
@@
    Finally, there’s the output problem, which is the converse of the input problem [18]. The challenge of encoding classical information into quantum states also extends to reading the output quantum state |$\ket{x}$ into classical information $\bm{x}$ when HHL is finished. Learning the value of any specific entry $x_{i}$ by measurement will, in general, require repeating the algorithm roughly $N$ times, negating the significant speedup.
To summarize, HHL is an algorithm designed to solve a system of linear equations approximately in logarithmic time. It serves as a template for other quantum algorithms, as long as one can address and mitigate the various caveats associated with its practical implementation [1]. If these challenges are properly handled, HHL can find applications in various real-world scenarios.
3.3.1 Large scale machine learning models in fault tolerant era
@@
The paper [90] introduces a technique known as Carleman linearization, demonstrating that non-linear ODEs can be linearized into an infinite set of linear ODEs that can be truncated to a chosen level of approximation. This allows the approximate solution of challenging non-linear ODEs by solving the easier linearized versions. The authors establish that if the original ODEs are non-linear and dissipative, then the error resulting from the truncation in Carleman linearization can be controlled. This implies that a non-linear, dissipative system can be efficiently linearized.
@@
Dissipativeness plays a crucial role in various machine learning tasks. Many machine learning processes, including gradient descent, exhibit dissipative behavior. Treating gradient descent as dynamics reveals that it operates as an open first-order system outside the scope of Lagrangian mechanics [89]. Biological processes in the brain also exemplify the significance of dissipation, where forgetting becomes essential for making room to remember new information [89].
The findings of [89] suggest the potential for an efficient algorithm to solve stochastic gradient descent. Generalizing the techniques for differential equations to ordinary difference equations, it is observed that in the learning process, the Hessian eigenvalues characterize the degree of dissipation. Towards the end of training, there are roughly equal numbers of positive and negative eigenvalues, signifying comparable dissipative and non-dissipative modes. However, at the beginning of training, the number of dissipative modes significantly exceeds the number of non-dissipative modes, creating highly asymmetric Hessian eigenvalues. This phase is recognized as highly dissipative, consistent with the intuition that the system learns rapidly in the initial stages. The ODE solver described earlier is applicable during this phase, enabling the solution of machine learning tasks. Towards the end of training, when positive and negative eigenvalues become comparable, the algorithm may not be suitable for solving ODEs in general. An exception is when approaching a local minimum, where the Hessian matrix is positive definite, indicating a dissipative system suitable for the ODE solver. In summary, this quantum algorithm might help machine learning processes significantly, and if the Hessian becomes symmetric towards the end, classical algorithms can be employed for further computation.
@@
Indeed, the vision of leveraging quantum algorithms, particularly with large-scale implementations of algorithms like the HHL, holds the promise of addressing the increasing costs associated with classical algorithms. The potential gains in efficiency and computational speed offered by quantum computers could revolutionize machine learning and other computational tasks. However, it’s important to emphasize that realizing these advancements requires the implementation of full quantum error correction to ensure fault tolerance. Overcoming the challenges associated with error correction remains a critical aspect of bringing quantum computing applications to fruition. Moreover, some further explorations about machine learning foundations should be given along this line, like improving the algorithms from solving gradient descent to solving the matrix multiplication problems in large-scale classical neural networks for back propagation, where quantum linear algebra techniques or parallel computing from analog photonic devices might be helpful for those tasks [14].
3.4 Quantum random access memory
@@
Now, let’s explore the challenges associated with the “input problem” and the “output problem”. Classical data needs to be inputted into a quantum computer before undergoing processing [18]. Referred to as the ‘input problem’, this step is generally performed with minimal overhead but can pose a significant bottleneck for certain algorithms. Similarly, the ‘output problem’ arises when retrieving data after it has been processed on a quantum device. Much like the input problem, the output problem can lead to a substantial operational slowdown.
@@
In particular, when considering the application of algorithms like HHL, least squares fitting, quantum support vector machines, and related approaches, substantial amounts of classical data might be loaded into a quantum system. This loading process can demand significant time, as acknowledged in the literature [1, 18]. While the use of a quantum random access memory (QRAM) [54, 56] might, in principle, address this issue, the associated costs might be prohibitive for large-scale data problems [10].
@@
QRAM can be conceptualized as an architecture designed for the implementation of quantum oracles [54]. Consider a computational problem where the input is represented by a classical data vector $\bm{x}\leavevmode\nobreak\ .$ An oracle, often described as a black box, is a tool that can be interrogated to disclose information about $\bm{x}\leavevmode\nobreak\ .$ Although the oracle reveals details about $\bm{x}$ upon request, the specific method it employs to obtain this information is not explicitly specified.
Figure 7: Classical and quantum data-lookup oracles [54].
@@
Oracles can exist in either classical or quantum forms, as illustrated in FIG. 7. A basic example of a classical oracle is the data-lookup oracle. This type of oracle is activated by providing it with an index $i$ as input, and it subsequently outputs the corresponding vector element $x_{i}\leavevmode\nobreak\ .$ A natural extension of this classical data-lookup oracle is the quantum data-lookup oracle. In the quantum scenario, both the inputs and outputs of the oracle query are represented as quantum states. The query itself is implemented by a unitary operation, denoted as $\mathcal{O}^{(\text{DL})}_{x}$, which performs the mapping [54]:
@@
where the notation $b$ represents any computational basis state and $\oplus$ denotes addition modulo 2. The superscripts $A$ and $B$ refer to two quantum registers, the state of register $A$ indicates which element to look up, and the query encodes this element into the state of register $B\leavevmode\nobreak\ .$ It’s important to note that, while the inputs and outputs of the query are quantum, the data being queried is classical. This characteristic makes quantum oracles serve as a bridge between classical data and quantum algorithms [54].
However, merely loading classical data into quantum memory is not sufficient; the process needs to be fast [1]. QRAM is designed to address this requirement. In QRAM, the quantum state summarizing the vector uses $\log{d}$ qubits, and the QRAM operation involves $\mathcal{O}(d)$ operations distributed over $\mathcal{O}(\log{d})$ steps, which can be performed in parallel [18].
For QRAM, a quantum superposition of different addresses |$\ket{\psi_{\text{in}}}$ serves as input, and the QRAM produces an entangled state |$\ket{\psi_{\text{out}}}$ where each address is correlated with the corresponding memory element [54]:
@@
Figure 8: The left figure is a picturesque illustration of fanout QRAM [54] to retrieve information stored at $i=5\leavevmode\nobreak\ .$ On the other hand, the right figure is a picturesque illustration of bucket-brigade QRAM [54] to retrieve information stored at $i=5\leavevmode\nobreak\ .$
@@
The bucket-brigade QRAM is a bit different in the first stage. We don’t flip all the router qubits in the layer corresponding to the address qubits. Instead, we route the address qubits into the QRAM one by one, with the router qubits initialized at some state |$\ket{w}\leavevmode\nobreak\ .$ Since we route in the address qubits in order, we naturally carve out the path that will lead to the information stored in $x_{5}\leavevmode\nobreak\ .$
@@
There are debates about whether QRAM is actually practical since it is time-saving but also hardware-consuming. Recently, there are several different QRAM designs that are hardware-efficient. One example is developed by the authors of [55]. They made use of circuit quantum acoustodynamic (cQAD) systems [103, 51] and constructed quantum gates by applying external drives with different frequencies.
@@
Furthermore, due to its logarithmic depth nature, when QRAM is stored with a large number of data, the speed for querying it will ultimately be bounded by the speed of light. The problem is how many qubits QRAM can store before we hit the boundary set by locality. This problem is addressed in [129], where the authors considered locality and found that some 2D or 3D QRAM designs can handle up to $10^{20}$ qubits before hitting the constraint line. This might be already a sufficient number of qubits to perform machine learning training programs. As an result, QRAM might be still reasonable to build in the future.
@@
Although it looks like that QRAM might be implementable, it is also important that there is no definitive evidence or proof in the hardware especially in the large scale up to date. So whether we can achieve exponential speed ups that some quantum machine learning programs promises might be still unknown and controversial. On the other hand, for some of the QML programs, one can go around QRAM by implementing a quantum subroutine which prepares the quantum states that encodes classical information [18].
4 Statistical learning theory
@@
In this section, our aim is to approach a quantum counterpart to statistical learning theory, initiating a systematic framework for delving into the fundamentals of quantum machine learning. A crucial aspect intertwined with statistical learning theory is shadow tomography, addressing a challenge rooted in the distinctive nature of quantum measurements—destructiveness. When concluding a Quantum Machine Learning (QML) program and obtaining a quantum state, extracting classical information becomes a non-trivial endeavor. The challenge arises from the inevitable collapse of the state during measurement, limiting access to the original state. Working with a single copy of an unknown quantum state, denoted as $\rho$, poses an inherent difficulty, as clever strategies cannot approximate a classical description of $\rho$ through direct measurement. The overarching task of reconstructing a description of a $D$-dimensional quantum mixed state $\rho$, given multiple copies of $\rho$, falls under the domain of quantum state tomography [2]. On the other hand, works such as [21, 22] studied statistical complexity in the context of quantum machine learning, which also closely relates to this topic.
@@
One significant development inspired by shadow tomography is known as classical shadow. As discussed earlier in section 2.1.3, this approach holds promise for addressing the output or downloading problem. Classical shadow infers the output state $\rho$ by performing random measurements, averaging the obtained classical data, and then inverting the results to retrieve the quantum state. This method extensively leverages basics of probability theory to control errors, making it an efficient quantum-to-classical data converter [41]. Utilizing classical shadow, one can devise classical machine learning algorithms that exhibit only polynomial differences with full fault-tolerant QML programs, particularly when focusing on average error [66]. Moreover, this approach holds significant applications in QML for quantum data or quantum simulators, especially when a readout is required at the end.
4.1 Shadow tomography